OpenStudy (jkbo):

please solve 9x^2 - 5 =29

OpenStudy (anonymous):

Try using the quadratic formula.

OpenStudy (anonymous):

Wait. Just add 5 to the other side and take the square root.

OpenStudy (anonymous):

9x^2=24

OpenStudy (anonymous):

and divide by 9

OpenStudy (anonymous):

x^2=24/9

OpenStudy (anonymous):

Then take the square root of both sides.

OpenStudy (anonymous):

Oh, sorry it would be 9x^2=34

OpenStudy (anonymous):

Well, just divide 34 by 9 then x^2=34/9 x=+/- sqrt(34/9)

OpenStudy (anonymous):

But since the 9 is a perfect root, you can write it as x=+/- sqrt(34)/3

OpenStudy (anonymous):

x^2=sqrt11.333333333333333333333333

OpenStudy (anonymous):

*Note: There are 2 answers. 1 positive, and 1 negative.

OpenStudy (jkbo):

yeah thats where I am suck you guys. at 9 divide by 34

OpenStudy (anonymous):



OpenStudy (jkbo):

ok I got -11.3333333

OpenStudy (jkbo):

@zordoloom I dont know what to do from there

OpenStudy (anonymous):

-11.33 is not correct. I posted the correct answers in the drawing.

OpenStudy (jkbo):

oo ok thank you very much